OpenStudy (lakers039):

1/"5 root 2" plus "one"

jhonyy9 (jhonyy9):

how is this more exactly please ?

jhonyy9 (jhonyy9):

5 from there is the index of radical ? or what ?

jhonyy9 (jhonyy9):

\[1/(\sqrt[5]{2} +1)\]

jhonyy9 (jhonyy9):

hence ?

OpenStudy (kailey7277):

i think it's \[\frac{ 1 }{ 5\sqrt{2} }+1\]

OpenStudy (mathmale):

@lakers039: If other users have to question you about what you meant by 1/"5 root 2" plus "one" that's an indication that you need to provide the instructions that came with this problem. Would you do that now, please?

OpenStudy (mathmale):

I highly recommend that you type out this problem in math symbols, not in words.

OpenStudy (mathmale):

I'd suggest you re-post this problem. If you choose to do so, please be sure to include the instructions with your post.